# mybibliography.bib

### v1

@@ Abstract
A common technique for ameliorating the computational costs of running large neural models is sparsification, or the removal of neural connections during training. Sparse models are capable of maintaining the high accuracy of state of the art models, while functioning at the cost of more parsimonious models. The structures which underlie sparse architectures are, however, poorly understood and not consistent between differently trained models and sparsification schemes. In this paper, we propose a new technique for sparsification of recurrent neural nets (RNNs), called moduli regularization, in combination with magnitude pruning. Moduli regularization leverages the dynamical system induced by the recurrent structure to induce a geometric relationship between neurons in the hidden state of the RNN. By making our regularizing term explicitly geometric, we provide the first, to our knowledge, a priori description of the desired sparse architecture of our neural net. We verify the effectiveness of our scheme for navigation and natural language processing RNNs. Navigation is a structurally geometric task, for which there are known moduli spaces, and we show that regularization can be used to reach 90% sparsity while maintaining model performance only when coefficients are chosen in accordance with a suitable moduli space. Natural language processing, however, has no known moduli space in which computations are performed. Nevertheless, we show that moduli regularization induces more stable recurrent neural nets with a variety of moduli regularizers, and achieves high fidelity models at 98% sparsity.
@@ 1 Introduction
Sparsification, or the excision of neural connections during training, is an important technique in the manufacture of computationally efficient deep learning models. Neural nets used across applications are heavily over-parameterized, and therefore both resilient to substantial sparsification and prone to over-fitting [frankle2018lottery, gale2019state]. Regularization and pruning techniques have therefore found great success in improving the quality of models trained, while simultaneously decreasing the computational costs of training and running large neural architectures.
@@
While model sparsification has been attempted using many different heuristics [NEURIPS2022_sparsity, NEURIPS2022_089b592c_controlledsparsity, louizos2017learning, narang2017exploring, singh2020woodfisher], the underlying sparse architectures are poorly understood, and vary greatly subject to small changes in both the sequence that training data is presented [frankle2020linear] and the initial randomization of weights [frankle2018lottery, gale2019state]. Frankle, et al. (2018) proposed the Lottery Ticket Hypothesis as a potential explanation of the instability of sparse models [frankle2018lottery]. The Lottery Ticket Hypothesis suggests that sparse models arise from fortunate weight initializations on small subnetworks. Neural nets are designed to be heavily over-parameterized because this increases the probability of fortunate weight initializations on subnetworks, i.e. that some subnetwork “wins the lottery."
@@
This hypothesis is intuitive and appealing. However, it leaves open the problem of describing sparse neural architectures. Retraining neural nets on the so-called “winning lottery tickets" with re-randomized weights fails to recover high-quality networks in both Resnets and transformers [gale2019state]. This presents fundamental obstacles to any project of structured ab initio sparsity, though it does not rule out that some sparse architectures may be more stable than others. Recurrent neural nets (RNNs), however, have underlying geometric properties which may allow for greater interpretability of sparse architectures: Sussillo and Barak (2013), for instance, show that the hidden state of various low-dimensional RNNs is concentrated along slow manifolds of a dynamical system [sussillo2013openingtheblackbox].
@@
In this paper, we provide evidence that the geometry of RNN hidden states can be parlayed into geometric structures underlying sparse RNNs. We do this by combining magnitude pruning, in which the lowest valued weights of the neural net are thresholded during training, with moduli regularization. Moduli regularization, which we detail in Section 2.2, is a topologically inspired regularization term applied to the hidden-state update matrix of the RNN. It is computed by embedding hidden-state neurons into a metric space $\mathcal{M}$, and differentially penalizing weights according to their distance on $\mathcal{M}$.
@@
This approach reflects the continuous attractor theory of RNNs [seung1997learning]. The continuous attractor theory posits that the hidden layer of an RNN isolates a low dimensional parameter space (the eponymous continuous attractor, or the moduli space) of stable states, and updates at each time correspond to small movements on this space. From this perspective, moduli regularization is a method of training RNNs which favors particular embeddings of the continuous attractor into the hidden state space: specifically, embeddings which are sparse in the neural basis. By specifying neural connections based on distance on a manifold, we perform the inverse to a common strategy for manifold detection in neuroscience, in which a manifold is discovered by linking neurons that simultaneously fire [gardner2022toroidal].
@@
We test the efficacy of our method in two recurrent architectures: a navigation RNN [sorscher2023unified] and a natural language processing RNN [pytorch_examples]. In each case, we show that the choice of a well-suited moduli space and inhibitor function create superior sparse models as opposed to comparable methods. We then randomize the model’s weights and retrain the neural net on the sparse architecture to test the stability of the learned sparse architecture, showing that moduli regularization not only provides superior sparse models, but also produces sparse architectures which are more stable than alternative methods. To the best of our knowledge, this is the first paper to provide evidence of structured a priori sparsity in any neural architecture.
Our contributions are the following:
  • We introduce moduli regularization for recurrent neural networks.
  • We show that regularization premised on well-chosen moduli spaces and inhibitor functions creates superior sparse neural networks suitable for single-shot training and magnitude-based pruning. We demonstrate the effectiveness of this approach on three benchmarks.
  • We provide evidence that a priori sparse structure can be used to train superior sparse neural nets.
@@ 1.1 Related work
Continuous attractors. Attractor dynamics in RNNs and their utility in image classification were first proposed by [seung1997learning]. Many recent papers discover continuous attractors by analyzing the hidden state of trained RNNs [elman1991distributed, schaeffer2023self, sorscher2023unified, sussillo2013openingtheblackbox]. Similar continuous attractors, and handmade networks that instantiate them, are common in the analysis of neural firing in the brain [burak2009accurate, gardner2022toroidal, guanella2007model]. Work to encode favorable properties in RNNs, such as translation equivariance [zhang2022translation], are also studied in hand-crafted neural nets. However, these studies lack methods to inform the training of general RNNs with their underlying attractors.
@@
Topology. The incorporation of parameterizing manifolds into neural network architectures has received some attention: architectures built around Grassmannians [DBLP:journals/corr/HuangWG16], symmetric positive-definite matrices [DBLP:journals/corr/Huang0LLSGC16, konstantinidis2023multimanifold], and the Klein bottle [topdeeplearning] have been proposed for various tasks. These approaches are built around a priori knowledge of data moduli spaces, which is then built into the architecture. Other authors have attempted to combine simplicial complexes learned from the data [osti_10303167] and from neural data recorded from mice [mitchell2024topological].
@@
Sparsity. Investigations of regularization in recent literature includes both methods focused on achievement of fixed sparsity targets [NEURIPS2022_sparsity, NEURIPS2022_089b592c_controlledsparsity, louizos2017learning, singh2020woodfisher] and methods which sparsify differently depending on the distribution of weights [narang2017exploring]. However, in large models, simple magnitude pruning remains the state of the art in sparsification [gale2019state]. Training ab initio sparse models remains a fundamental challenge: sparse models are less stable than dense ones, meaning that random weight initializations are less likely to converge to high fidelity models [frankle2018lottery]. Even with identical weight initializations, Frankle et al. [frankle2020linear] show that the stability of network training with respect to randomness in stochastic gradient descent was achieved only after a degree of training, at which point the corresponding sparsified networks agreed.
@@
The Lottery Ticket Hypothesis of [frankle2018lottery] is a proposed explanation for the paradox of the effectiveness of large-scale sparsification, and the ineffectiveness of training a priori sparse models [gale2019state]. It argues that the large over-parameterization of neural nets makes it very likely that some subnetwork is configured with a fortunate set of weights. Much past work on sparse neural nets focuses on convolutional layers of image recognition neural networks, where benchmarks are well established [gale2019state, louizos2017learning, molchanov2017variational, zhu2017prune], though attention has also been given to transformers [gale2019state, vaswani2017attention] and recurrent neural nets [narang2017exploring] for natural language processing.
@@ 2.1 Preliminaries
A recurrent neural net is a general term for any neural net designed to analyze sequences or streams of data, one element at a time. The Elman Recurrent Neural Network, henceforward abbreviated as RNN, is one of the most straightforward yet effective structures in common usage[elman1991distributed]. The RNN receives a stream of data $\{x_{0},x_{1},x_{2},...\}$ represented as vectors, and maintains a hidden state $H_{t}$ at each time $t$, a vector that functions as an internal representation of the state of the system. After an initial hidden state $H_{-1}$ is fixed, the $H_{t}$ are computed recursively as
@@
where $B_{d}$ is a vector, called the decoder bias, and $s^{\prime}$ is an activation function, such as the softmax (Equation (8)) for classification problems, or the identity matrix for regression problems.
An RNN can be viewed as a discrete approximation of a dynamical system on $\mathbb{R}^{n}$. It creates a vector field on $\mathbb{R}^{n}$, which is an assignment, for each point $x\in\mathbb{R}^{n}$, of a vector $v(x)\in\mathbb{R}^{n}$ such that $v(x)$ is smoothly varying in $x$. To flow along a vector field from a point $x_{0}$ means to take small steps in the direction $v(x_{t})$ along a sequence of points $x_{t}=x_{t-1}+v(x_{t-1})dt$. A fixed point $x$ of a vector field is a point with velocity $v(x)=0$. In addition, such a point $x$ is called stable if it flows from points in a neighborhood $U$ converge to $x\in U$. A continuous attractor associated to a vector field $\mathcal{A}$ is a continuous locus of stable points.
@@ 2.2 Moduli regularization for RNNs
$L_{1}$ regularization [tibshirani96, l1_paper] does not directly create sparse networks, and much work [gale2019state, louizos2017learning, molchanov2017variational, narang2017exploring, zhu2017prune] has gone into creating different regularization techniques that will more directly induce sparsity. Our method, which we term moduli regularization, is rooted in the hypothesis that manifolds (Appendix B) usefully approximate the hidden space of RNNs.
We fix a metric space $(\mathcal{M},d_{\mathcal{M}})$. Let $\{e_{1},...,e_{n}\}$ be the standard basis vectors for the hidden state $\mathbb{R}^{n}$ of an RNN, and choose an embedding $i:\{1,...,n\}\to\mathcal{M}$ from the set of hidden state neurons into $\mathcal{M}$. Let $W_{hh}=(w_{jk})_{1\leq j,k\leq n}$ be the hidden update matrix of Equation (1). For a function $f:\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ and $\ell\geq 1$, define an $f$-regularizer $R_{f}$ of $W_{hh}$ associated to the embedding $i$ to be
@@
Figure 1: (a) Diagrammatic structure of the Elman RNN. (b) Hidden state neurons, represented by blue dots, are embedded into a moduli space, the torus. (c) Sparsification of the hidden update matrix of an RNN (Equation 1). Above depicts random sparsification, and below depicts sparsification in line with moduli regularization (briefly, moduli sparsification). Yellow boxed points are neurons with a non-zero weight connecting them to the center neuron. Moduli sparsification respects the geometry of the chosen moduli space, which is ignored by standard sparsification techniques.
The function $f:\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ is referred to as the inhibitor function. The inhibitor function is an additional parameter of the regularizer, which dictates the manner in which information is passed along the metric space. For example, explicit constructions of continuous attractors in neural nets for navigation [burak2009accurate, guanella2007model] are constructed with connectivity given by the Ricker wavelet
@@
where $d$ is the distance between two points on the moduli space. The attractors are unstable if simpler functions, such as $f(d)=d$, are applied as inhibitor functions, suggesting that an apt inhibitor choice is important for moduli regularizers to function. We diagram the circle, sphere, torus, and Klein bottle with various inhibitor functions in Figure 2. Further discussion and results on inhibitor functions can be found in Appendix E, and a heatmap from a trained RNN can be found in Appendix F.
Figure 2: A neuron $y$ is fixed. Red points depict output neurons $z$ with low regularizing values, and therefore potentially large weights $w_{yz}$, while blue points are output neurons $z^{\prime}$ with high regularizing values, and therefore smaller potential weights $w_{yz^{\prime}}$. Row (1) depicts the circle, sphere, and Klein bottle, while row (2) depicts the torus. Column (a) has inhibitor function $f(d)=d^{2}$, column (b) has inhibitor function $f(d)=10-10(e^{-\frac{d^{2}}{5}}-e^{-d^{2}})$, and column (c) has inhibitor function $f(d)=10+10\cos(2d)$.
A multi-layer RNN is a stack of recurrent neural networks: the first recurrent layer takes as input the input vectors $x_{t}$, as in Equation (1). Subsequent recurrent layers take as input the previous layer’s hidden state $H_{t}$. We regularize each layer of multi-layer RNNs independently on the same manifold by choosing three different embeddings and using these to generate regularizers. It is also possible to change the manifold between recurrent layers, however we do not present any such experiments. Furthermore, this framework could be studied for non-recurrent architectures, although that is not the focus of the current work.
2.3 Choice of embeddings
The moduli regularizer involves the choice of an embedding $i:\{1,2,...,n\}\to\mathcal{M}$ in addition to the choice of $\mathcal{M}$. We initialize our embeddings by drawing points independently and uniformly at random over $\mathcal{M}$. When $\mathcal{M}$ has a smooth structure and the inhibitor function $f$ is differentiable, $i$ can be taken as a parameter of the RNN and optimized by gradient descent during the course of training.
Monotonically increasing inhibitor functions cannot be effectively used with trained embeddings: in this setting, the moduli regularizer (Equation (5)) is minimized by any embedding $\{j\}_{1\leq j\leq n}\to*\hookrightarrow\mathcal{M}$ which maps all neurons to a single point in the moduli space. Moduli regularization then degenerates to $L_{\ell}$ regularization with regularizing coefficient $f(0)$. Other choices of $f$, however, such as the difference of Gaussians

### v2

# Geometric sparsification in recurrent neural networks

@@ Abstract
A common technique for ameliorating the computational costs of running large neural models is sparsification, or the pruning of neural connections during training. Sparse models are capable of maintaining the high accuracy of state of the art models, while functioning at the cost of more parsimonious models. The structures which underlie sparse architectures are, however, poorly understood and not consistent between differently trained models and sparsification schemes. In this paper, we propose a new technique for sparsification of recurrent neural nets (RNNs), called moduli regularization, in combination with magnitude pruning. Moduli regularization leverages the dynamical system induced by the recurrent structure to induce a geometric relationship between neurons in the hidden state of the RNN. By making our regularizing term explicitly geometric, we provide the first, to our knowledge, a priori description of the desired sparse architecture of our neural net, as well as explicit end-to-end learning of RNN geometry. We verify the effectiveness of our scheme under diverse conditions, testing in navigation, natural language processing, and addition RNNs. Navigation is a structurally geometric task, for which there are known moduli spaces, and we show that regularization can be used to reach 90% sparsity while maintaining model performance only when coefficients are chosen in accordance with a suitable moduli space. Natural language processing and addition, however, have no known moduli space in which computations are performed. Nevertheless, we show that moduli regularization induces more stable recurrent neural nets, and achieves high fidelity models above 90% sparsity.
@@ 1 Introduction
Sparsification, or the excision of neural connections during training, is an important technique in the manufacture of computationally efficient deep learning models. Neural nets used across applications are heavily over-parameterized, and therefore both resilient to substantial sparsification and prone to over-fitting [9, 11]. Regularization and pruning techniques have therefore found great success in improving the quality of models trained, while simultaneously decreasing the computational costs of training and running large neural architectures.
@@
While model sparsification has been attempted using many different heuristics [12, 13, 30, 38, 47], the underlying sparse architectures are poorly understood, and vary greatly subject to small changes in both the sequence that training data is presented [10] and the initial randomization of weights [9, 11]. Frankle, et al. (2018) proposed the Lottery Ticket Hypothesis as a potential explanation of the instability of sparse models [9]. The Lottery Ticket Hypothesis suggests that sparse models arise from fortunate weight initializations on small subnetworks. Neural nets are designed to be heavily over-parameterized because this increases the probability of fortunate weight initializations on subnetworks, i.e. that some subnetwork “wins the lottery."
@@
This hypothesis is intuitive and appealing. However, it leaves open the problem of describing sparse neural architectures. Retraining neural nets on the so-called “winning lottery tickets" with re-randomized weights fails to recover high-quality networks in both Resnets and transformers [11]. This presents fundamental obstacles to any project of structured ab initio sparsity, though it does not rule out that some sparse architectures may be more stable than others. Recurrent neural nets (RNNs), however, have underlying geometric properties which may allow for greater interpretability of sparse architectures: Sussillo and Barak (2013), for instance, show that the hidden state of various low-dimensional RNNs is concentrated along slow manifolds of a dynamical system [50], also described as attractor manifolds.
@@
In this paper, we provide evidence that the geometry of RNN hidden states can be parlayed into geometric structures underlying sparse RNNs. We do this by combining magnitude pruning, in which the lowest valued weights of the neural net are thresholded during training, with moduli regularization. Moduli regularization, which we introduce in Section 2.2, is a topologically inspired regularization term applied to the hidden-state update matrix of the RNN. It is computed by embedding hidden-state neurons into a metric space $\mathcal{M}$, and differentially penalizing weights according to their distance on $\mathcal{M}$. We further show that the geometry of $\mathcal{M}$ can be implicitly learned and implemented during training at relatively low cost.
@@
Our approach reflects the continuous attractor theory of RNNs [45]. The continuous attractor theory posits that the hidden layer of an RNN isolates a low dimensional parameter space (the eponymous continuous attractor, or the moduli space) of stable states, and updates at each time correspond to small movements on this space. From this perspective, moduli regularization is a method of training RNNs which favors particular embeddings of the continuous attractor into the hidden state space: specifically, embeddings which are sparse in the neural basis. By specifying neural connections based on distance on a manifold, we perform the inverse to a common strategy for manifold detection in neuroscience, in which a manifold is discovered by linking neurons that simultaneously fire [14].
@@
We test the efficacy of our method in three recurrent architectures: a navigation RNN [4], a natural language processing RNN [3], and an adding problem RNN [27]. In each case, we show that a moduli space well-suited to the RNN creates superior sparse models as opposed to comparable methods. We then randomize the model’s weights and retrain the neural net on the sparse architecture to test the stability of the learned sparse architecture, showing that moduli regularization not only provides superior sparse models, but also produces sparse architectures which are more stable than alternative methods.
To the best of our knowledge, this work is the first to demonstrate the role of structured a priori sparsity in RNNs. It further provides new evidence that global topological structures, rather than purely local network properties, play a critical role in determining the performance and stability of sparse neural architectures.
The key contributions of this work are:
  • We introduce moduli regularization as a framework for highly efficiently sparsifying RNNs using geometric principles.
  • We introduce the first method for end-to-end explicit learning of underlying geometric structures during RNN training.
  • We provide empirical validation of our method, highlighting the importance of global topological structure, as opposed to purely local network structures, in sparse network design.
Our approach advances a structured and principled methodology for sparsity in neural networks, addressing both theoretical and practical aspects of efficient and robust model design.
@@ 1.1 Related work
Continuous attractors. Attractor dynamics in RNNs and their utility in image classification were first proposed by [45]. Many recent papers discover continuous attractors by analyzing the hidden state of trained RNNs [6, 43, 49, 50]. Similar continuous attractors, and handmade networks that instantiate them, are common in the analysis of neural firing in the brain [2, 14, 18]. Work to encode favorable properties in RNNs, such as translation equivariance [56], are also studied in hand-crafted neural nets. However, these studies lack methods to inform the training of general RNNs with their underlying attractors.
@@
Topology. The incorporation of parameterizing manifolds into neural network architectures has received some attention: architectures built around Grassmannians [24], symmetric positive-definite matrices [23, 26], and the Klein bottle [31] have been proposed for various tasks. These approaches are built around a priori knowledge of data moduli spaces, which is then built into the architecture. Other authors have attempted to combine simplicial complexes learned from the data [40] and from neural data recorded from mice [35].
Manifold discovery. A foundational challenge for topologically aware neural nets is manifold discovery, or the learning of low dimensional manifolds that well describe given data [7, 51]. There is growing interest in the topological data and machine learning communities regarding the use of mixed curvature representations as a replacement for standard Euclidean representations of both data and learned features. This allows representations that flexibly match the true curvature of the dataset by learning both optimal curvature values for each product component [17, 48] and the distribution of curvature signs throughout the product [42]. These techniques find notable improvements over baseline representations in variational autoencoders [42], embedding spaces [17], and learned representations for downstream tasks [42].
@@
Sparsity. Investigations of regularization in recent literature includes both methods focused on achievement of fixed sparsity targets [12, 13, 30, 47] and methods which sparsify differently depending on the distribution of weights [38]. However, in large models, simple magnitude pruning remains the state of the art in sparsification [11]. Training ab initio sparse models remains a fundamental challenge: sparse models are less stable than dense ones, meaning that random weight initializations are less likely to converge to high fidelity models [9]. Even with identical weight initializations, Frankle et al. [10] show that the stability of network training with respect to randomness in stochastic gradient descent was achieved only after a degree of training, at which point the corresponding sparsified networks agreed.
@@
The Lottery Ticket Hypothesis of [9] is a proposed explanation for the paradox of the effectiveness of large-scale sparsification, and the ineffectiveness of training a priori sparse models [11]. It argues that the large over-parameterization of neural nets makes it very likely that some subnetwork is configured with a fortunate set of weights. Much past work on sparse neural nets focuses on convolutional layers of image recognition neural networks, where benchmarks are well established [11, 30, 36, 57], though attention has also been given to transformers [11, 54] and recurrent neural nets [38] for natural language processing.
@@ 2.1 Preliminaries
A recurrent neural net is a general term for any neural net designed to analyze sequences or streams of data, one element at a time. The Elman Recurrent Neural Network, henceforward abbreviated as RNN, is one of the most straightforward yet effective structures in common usage [6]. The RNN receives a stream of data $\{x_{0},x_{1},x_{2},...\}$ represented as vectors, and maintains a hidden state $H_{t}$ at each time $t$, a vector that functions as an internal representation of the state of the system. After an initial hidden state $H_{-1}$ is fixed, the $H_{t}$ are computed recursively as
@@
where $B_{d}$ is a vector, called the decoder bias, and $s^{\prime}$ is an activation function, such as the softmax (Equation (9)). A multi-layer RNN is a stack of recurrent neural networks: the first recurrent layer takes as input the input vectors $x_{t}$, as in Equation (1). Subsequent recurrent layers take as input the previous layer’s hidden state $H_{t}$. For an $n$-layer RNN, with fixed initial hidden states $\{H_{-1,k}\}_{k=0,...,n-1}$, the hidden states are updated recursively as
$H_{t,k}=s(W_{hh}^{k}H_{t-1,k}+W_{ih}H_{t,k-1}+B),$ (3)
where $s,W,H,B$ are as above when $k>0$, and using the update of (1) when $k=0$.
An RNN can be viewed as a discrete approximation of a dynamical system on $\mathbb{R}^{n}$. The RNN acts as a continuous automorphism of $\mathbb{R}^{n}$, which induces a vector field on $\mathbb{R}^{n}$. A vector field is an assignment, for each point $x\in\mathbb{R}^{n}$, of a vector $v(x)\in\mathbb{R}^{n}$ such that $v(x)$ is smoothly varying in $x$. To flow along a vector field from a point $x_{0}$ means to take small steps in the direction $v(x_{t})$ along a sequence of points $x_{t}=x_{t-1}+v(x_{t-1})dt$. A fixed point $x$ of a vector field is a point with velocity $v(x)=0$. In addition, such a point $x$ is called stable if it flows from points in a neighborhood $U$ converge to $x\in U$. A continuous attractor associated to a vector field $\mathcal{A}$ is a continuous locus of stable points.
Continuous attractors have a metric structure, which means they are equipped with a distance function which satisfies the triangle inequality, is symmetric, and satisfies $d(x,y)=0$ if and only if $x=y$. A smooth manifold is a special kind of topological space which is locally homeomorphic to $\mathbb{R}^{n}$ (see Appendix B for details), and on which derivatives can be defined. Smooth manifolds are metric spaces whose distance function is called geodesic distance. Geodesic distance measures the length of the shortest path between two given points. Common examples of manifolds include a circle, sphere, torus, and $\mathbb{R}^{n}$.
@@ 2.2 Moduli regularization for RNNs
$L_{1}$ regularization [1, 44] does not directly create sparse networks, and much work [11, 30, 36, 38, 57] has gone into creating different regularization techniques that will more directly induce sparsity. Our method, which we term moduli regularization, is rooted in the hypothesis that manifolds (Appendix B) usefully approximate the hidden space of RNNs. This is a form of the manifold hypothesis, which suggests that high dimensional data is often encoded in a low dimensional manifold [8]. The manifold hypothesis has led to a wide variety of techniques for manifold learning and machine learning (e.g. [51, 7, 42, 48]). In RNNs the manifold hypothesis has been studied primarily in terms of manifold emergence during learning, such as the slow manifolds of [50] for state-change tasks, or toroidal grid cell emergence [4, 14, 49, 43]. Moduli regularization proposes immediate exploitation of the manifold structure to create superior sparse RNNs.
Fix a metric space $(\mathcal{M},d_{\mathcal{M}})$. Let $\{e_{1},...,e_{n}\}$ be the standard basis vectors for the hidden state $\mathbb{R}^{n}$ of an RNN, and choose an embedding $i:\{1,...,n\}\to\mathcal{M}$ from the set of hidden state neurons into $\mathcal{M}$. Let $W_{hh}=(w_{jk})_{1\leq j,k\leq n}$ be the hidden update matrix of Equation (1). For a function $f:\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ and $\ell\geq 1$, define an $f$-regularizer $R_{f}:\mathbb{R}^{n^{2}}\to\mathbb{R}$ of $W_{hh}$ associated to the embedding $i$ to be
@@
Figure 1: (a) Top: a traditional RNN tracks hidden states as a sequence of vectors in $\mathbb{R}^{n}$, without fixed structure; each bar represents the number stored at that dimension in the RNN. Bottom: with moduli regularization, the hidden state approximates a function on the circle, by using neurons (marked by dark lines) as a discrete approximation of the circle. These neurons record values approximating a distribution, represented by the height of the dark lines. (b) Hidden state neurons, represented by blue dots, are embedded into a moduli space, the torus. (c) Sparsification of the hidden update matrix of an RNN (Equations 1, 3). Above depicts random sparsification, and below depicts sparsification in line with moduli regularization (briefly, moduli sparsification). Yellow boxed points are neurons with a non-zero weight connecting them to the center neuron. Moduli sparsification respects the geometry of the chosen moduli space, which is ignored by standard sparsification techniques.
The function $f:\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ is referred to as the inhibitor function. The inhibitor function is an additional parameter of the regularizer, which dictates the manner in which information is passed along the metric space. For example, explicit constructions of continuous attractors in neural nets for navigation [2, 18] are constructed with connectivity given by the Ricker wavelet
@@
where $d$ is the distance between two points on the moduli space.
We regularize each layer of multi-layer RNNs independently on the same manifold by choosing three different embeddings and using these to generate regularizers. It is also possible to change the manifold between recurrent layers, however we do not present any such experiments. Furthermore, this framework could be studied for non-recurrent architectures, although that is not the focus of the current work (see Appendix I).
2.3 Learning the moduli space
The moduli regularizer involves the choice of an embedding $i:\{1,2,...,n\}\to\mathcal{M}$ in addition to the choice of $\mathcal{M}$. We initialize our embeddings by drawing points independently and uniformly at random over $\mathcal{M}$ when $\mathcal{M}$ is compact. When $\mathcal{M}$ has a smooth structure and the inhibitor function $f$ is differentiable, $i$ can be taken as a parameter of the RNN and optimized by gradient descent during the course of training. By choosing $\mathcal{M}$ to be Euclidean space $\mathbb{R}^{k}$, the embeddings can cluster on a submanifold, explicitly learning desirable geometric relationships—moreover, any manifold of dimensions less than or equal to $k/2$ can be perfectly approximated by this process, by the Whitney embedding theorem.
Monotonically increasing inhibitor functions cannot be effectively combined with trained embeddings: if used, the moduli regularizer (Equation (6)) is minimized by any embedding $\{j\}_{1\leq j\leq n}\to*\hookrightarrow\mathcal{M}$ which maps all neurons to a single point in the moduli space. Moduli regularization then degenerates to $L_{\ell}$ regularization with regularizing coefficient $f(0)$. Other choices of $f$, however, such as the difference of Gaussians